a man has a son and daughter. the man is four times as old as his daughter who is two years younger than her brother. in five years, the man will be three times as old as his son. how old is the man now?
Answer by ankor@dixie-net.com(22740)

a man has a son and daughter. the man is four times as old as his daughter who
is two years younger than her brother. in five years, the man will be three
times as old as his son. how old is the man now?
:
Let m = mans age
Let d = daughter's age
let s = son's age
:
write equations for each given statement:
;
"the man is four times as old as his daughter who is two years younger than her brother."
m = 4d
d = s-2
therefore:
m = 4(s-2)
m = 4s - 8
:
"in five years, the man will be three times as old as his son."
m + 5 = 3(s+5)
m + 5 = 3s + 15
m = 3s + 15 - 5
m = 3s + 10
Substitute (4s-8) for m, and find s
4s - 8 = 3s + 10
4s - 3s = 10 + 8
s = 18 yrs is the son's present age
:
how old is the man now?
:
m = 4s - 8
m = 4(18) - 8
m = 72 - 8
m = 64 yrs old now
:
:
Check solution in the statement:
in five years, the man will be three times as old as his son.
64 + 5 = 3(18+5)
69 = 3(23)